# Who Connects Global Aid? The Hidden Geometry of 10 Million Transactions

### v1

@@ Introduction
The modern architecture of global aid was born in the aftermath of World War II. In its infancy, Global Aid was conceived as a state-led instrument of geopolitical reconstruction managed by the OECD Development Assistance Committee[fuhrer1994story]. It was driven primarily by diplomatic and domestic political imperatives, with the allocation often determined by donor strategic interests rather than recipient need [dudley1976model, alesina2000gives, lancaster2007foreign]. However, the late 20th and early 21st centuries witnessed a radical phase transition[severino2009end]. A triple revolution characterised by new goals, new instruments and an explosion of new actors has reshaped the landscape[severino2010end]. This includes the rise of non-DAC donors[kharas2009new], the proliferation of vertical funds and private foundations, and a shift toward the complex interdependence of the Sustainable Development Goals[lee2016transforming].
Despite the growing scale, the politics of measurement remain contentious[weaver2010politics]. High-level policy attempts to harmonise this sprawling system, such as the Paris Declaration[oecd2005paris], have largely failed. Empirical analysis suggests that coordination efforts have not reduced proliferation[aldasoro2010less]. This leaves the sector in a state of unmanaged complexity, where uncoordinated high volumes of aid can sometimes hinder economic growth[djankov2008curse].
@@ Results
Recent network literature challenges the traditional volume hypothesis[burnside2000aid, collier2002aid, sachs2008end], suggesting instead that a recipient’s integration into the global network is a stronger predictor of growth and treaty ratification than raw financial flows[swiss2016world, swiss2017foreign].
@@
Figure 2: The hidden geometry of the aid ecosystem. UMAP dimensionality reduction projects high-dimensional network embeddings into 2D space to reveal functional clusters defined by two primary axes. The horizontal axis separates Humanitarian actors (left) from long-term Development institutions (right). The vertical axis differentiates Funders (top) from Implementers (bottom). This orthogonal arrangement creates four distinct quadrants populated by exemplars such as OCHA (Humanitarian Funder), Save the Children (Humanitarian Implementer), UNDP (Development Funder) and Chemonics (Development Implementer). The analysis also highlights the strategic position of connectors like J-PAL and The Hewlett Foundation which sit near the origin and bridge these functional zones.
@@
The separation between these clusters highlights a systemic coordination risk. The structural holes between the quadrants are bridged by relatively few actors, potentially stalling the transition from crisis response to resilience building. We identify specific connectors such as J-PAL and The Hewlett Foundation that sit near the origin of this coordinate system, suggesting they are uniquely capable of integrating these otherwise disconnected functional zones. Similar topological fissures have been observed in global health networks[shiffman2016emergence] and climate adaptation finance[weiler2021donor], where issue-specific clusters often fail to integrate with broader systems strengthening[han2018social, fergus2022power].
@@
To derive this ranking, we analysed the underlying bipartite graph connecting providers to receivers. We calculated the Network Centrality (Hub Score [kleinberg1999authoritative]) for all organisations to measure their systemic influence. In this bipartite context, a high Hub Score identifies actors that serve as critical “hubs”, connecting to a wide and significant range of recipients across the global network.
@@
We identify a core inner ring of 25 actors who dominate the network’s structural connectivity. This visualisation reveals a critical paradox between scale and influence. While the inner ring includes the expected financial giants (US Department of State, European Commission, World Bank) it also contains a distinct class of Knowledge Brokers. These are organisations with lower financial volume but disproportionately high network centrality (Extended Data Fig. LABEL:fig:ext_fig3).
@@
Their high betweenness centrality [anthonisse1971rush] (Extended Data Fig.3) indicates both J-Pal and the Hewlett foundation sit on the shortest paths between otherwise poorly connected donor-implementer clusters. These findings overturn the assumption that influence is solely a function of budget. In the hidden geometry of global aid, smaller research-led organisations provide the essential connective tissue that holds the fragmented system together.
@@ Discussion
Structural Fragmentation. Our topological analysis reveals that the global aid system is not a cohesive or integrated network. It is rather a collection of functional and geographic islands that are sparsely connected. The clear separation observed in the UMAP embeddings between the humanitarian and development clusters points to a persistent structural divide. Organisations tend to cluster with peers that share similar operational timeframes and funding sources[maoz2010networks, kinne2013network]. While this specialisation allows for operational efficiency within clusters it creates systemic risk at the macro level. It enables rapid response for humanitarian actors or long-term infrastructure planning for development banks but leaves the wider system disjointed.
@@
The sparseness of connections between these clusters suggests a high potential for duplication of effort and coordination failure. The economic costs of such fragmentation are well documented, often overwhelming the bureaucratic capacity of recipient nations[acharya2006proliferation, knack2007donor] and creating game-theoretic equilibria where donors prioritise visibility over collective impact[torsvik2005foreign]. When evidence is generated in one cluster such as a randomised control trial on health interventions within the development cluster the lack of structural bridges prevents it from diffusing efficiently to the humanitarian cluster. This fragmentation explains a longstanding paradox in development economics regarding why evidence of what works often fails to scale. The network geometry confirms that the barriers to scaling are not merely financial or political[findley2011localized, kinne2013network] but topological. Information simply cannot traverse the structural holes between the islands of the aid system without the intervention of specific bridging actors[burt1992social].
@@
The Strategic Role of Brokers. The identification of a solar system architecture offers a new strategic framework for donors. Traditionally aid portfolios are constructed based on volume. Funding is allocated to doers such as large NGOs and contractors capable of absorbing significant capital. However our centrality analysis demonstrates that high-volume implementers often operate within deep but narrow silos. Industry analysis of donor practices confirms that this fragmentation creates a “vicious cycle” of grant dependence where short-term reporting requirements stifle the long-term systemic growth of support organisations[Tompson2023sustain].
@@
To de-risk portfolios and accelerate systemic coordination donors must balance investments in doers with investments in connectors. Our results highlight that knowledge brokers provide disproportionate structural influence relative to their size. These are specifically academic research centers like the Abdul Latif Jameel Poverty Action Lab (J-PAL) at MIT and universities like Johns Hopkins University as well as strategic foundations like The Hewlett Foundation. These actors possess high betweenness centrality which means they sit on the shortest paths between otherwise disconnected nodes. Sociological theory posits that such weak ties are more effective for the diffusion of new information than strong ties to established partners[granovetter1973strength]. This aligns with the rise of “knowledge actors” who exert influence through transnational standard-setting rather than financial weight[stone2013knowledge, littoz2024knowledge, stone2010private]. J-PAL, for instance, leveraged this position to shift the sector toward evidence-based micro-interventions[karlan2011more, duflo2006field, glennerster2011small].
@@
Network Resilience. From a complex systems perspective the current architecture of global aid exhibits signs of structural fragility. The system is heavily dependent on the Donor Core. This consists of a small number of massive hubs like USAID, the World Bank and the European Commission that provide the majority of financial fuel. In network theory systems dependent on a few central hubs are resilient to random failure but highly vulnerable to targeted disruption or the abrupt removal of a hub[albert2000error, haldane2011systemic]. This vulnerability is not hypothetical. Recent volatility such as the recalibration of US multilateral commitments between 2017 and 2021 demonstrates how policy discontinuities in a key donor nation can rapidly propagate instability across the wider network.
Strengthening the middle layer of knowledge brokers creates a more resilient and adaptive mesh. Much like the internet’s distributed architecture allows data to reroute around damaged nodes—unlike a centralised telephone exchange which fails if the switchboard goes down, a network with a strong middle layer of brokers is less prone to fragmentation even if a major donor pulls back. These brokers act as the system’s memory and nervous system by retaining evidence and maintaining relationships that transcend political cycles. Our analysis suggests that the resilience of the global aid architecture depends on shifting from a star topology of centralised control to a mesh topology of distributed coordination[ostrom2017polycentric]. The hidden geometry we identified suggests that this mesh already exists in embryonic form through the university and foundation sectors. It simply requires recognition and resource support to function as a true stabiliser. Recent systemic design scholarship reinforces this arguing that traditional management dimensions must shift toward a “complexity paradigm” to handle such emergent dynamics[sobel2025complexity].
@@ Methods
We derived the global aid network from the International Aid Transparency Initiative (IATI) registry[fergus2022power]. This is the largest open-data standard for development cooperation. The dataset covers the period from 1967 to 2025 comprising over 10 million individual transactions. We extracted the complete corpus of transaction records involving 2,456 publishing organisations. To ensure consistency we filtered for nonzero financial transactions and normalised organisation names using fuzzy string matching to resolve variations in naming conventions. For example we matched “USAID” with “United States Agency for International Development”.
@@
We modelled the aid ecosystem as a bipartite or two-mode graph consisting of Provider nodes and Receiver nodes. An edge connects a Provider and a Receiver if a financial transaction occurred between them[newman2003structure, horowitz2021rethinking, coscia2013structure]. To analyse the structural relationships between donors and implementers we performed a bipartite projection to create a Provider-Provider co-investment graph. In this projected network two organisations are connected if they fund or implement projects in the same receiver country or sector. Edges were weighted by the frequency of co-occurrence to emphasise relational strength over pure financial volume.
@@
To reveal the hidden geometry of the system and identify functional clusters we utilised Uniform Manifold Approximation and Projection (UMAP)[McInnes2018]. We first generated 100-dimensional node embeddings using the node2vec algorithm[grover2016node2vec] which captures the structural equivalence of actors within the network. These high-dimensional embeddings were then projected into a 2D manifold using UMAP. We utilised standard parameters including 15 neighbours and a minimum distance of 0.1 with a cosine metric. This approach separates actors not by geography but by their topological role. It groups organisations that share similar funding patterns and partnership profiles.
@@
We calculated network centrality metrics to identify the core architecture of the system. Degree Centrality was used to measure the total connectivity of actors. Betweenness Centrality was calculated to quantify brokerage[boldi2014axioms]. This identifies actors that lie on the shortest paths between other pairs of organisations[borgatti2005centrality]. We validated these offline network metrics by correlating them with online visibility metrics using network centrality measures of linkage patterns from the organisations’ websites. We found a strong positive Pearson correlation of 0.48 between the structural position of an organisation in the IATI network and the centrality of its website in the World Wide Web (Global PageRank) (Extended Data Fig.4). The solar system visualisation ranks the top 100 organisations by a composite centrality score with the inner ring representing the top 25 actors. To handle the scale of the graph we employed advanced approximation algorithms for neighbourhood functions[boldi2011hyperanf]. This methodological approach aligns with previous efforts to map aid flows[easterly2008does, lim2008structure, charles2022understanding] but extends the analysis to a global scale, allowing for the detection of critical transitions[scheffer2012anticipating] and systemic risks[helbing2013globally, hidalgo2009building].
@@
The dataset analysed during the current study is publicly available in the International Aid Transparency Initiative (IATI) Registry. The raw XML data can be accessed at [Insert IATI URL]. The processed network files and edge lists used for the topological analysis are available in the [Insert Name] repository at [Insert GitHub/Zenodo URL].
@@ Acknowledgements
<to be added>
@@
<to be added>

### v2

@@ Introduction
The modern architecture of global aid was born in the aftermath of World War II. In its infancy, Global Aid was conceived as a state-led instrument of geopolitical reconstruction managed by the OECD Development Assistance Committee[1]. It was driven primarily by diplomatic and domestic political imperatives, with the allocation often determined by donor strategic interests rather than recipient need [2, 3, 4]. However, the late 20th and early 21st centuries witnessed a radical phase transition[5]. A triple revolution characterised by new goals, new instruments and an explosion of new actors has reshaped the landscape[5]. This includes the rise of non-DAC donors[6], the proliferation of vertical funds and private foundations, and a shift toward the complex interdependence of the Sustainable Development Goals[7].
Despite the growing scale, the politics of measurement remain contentious[8]. High-level policy attempts to harmonise this sprawling system, such as the Paris Declaration[9], have largely failed. Empirical analysis suggests that coordination efforts have not reduced proliferation[10]. This leaves the sector in a state of unmanaged complexity, where uncoordinated high volumes of aid can sometimes hinder economic growth[11].
@@ Results
Recent network literature challenges the traditional volume hypothesis[12, 13, 14], suggesting instead that a recipient’s integration into the global network is a stronger predictor of growth and treaty ratification than raw financial flows[15, 16].
@@
Figure 2: The hidden geometry of the aid ecosystem. UMAP dimensionality reduction projects high-dimensional network embeddings into 2D space to reveal functional clusters defined by two primary axes of network behaviour. The horizontal axis represents a network tendency spectrum, separating actors with Humanitarian-like transaction patterns (left) from those with Development-like patterns (right). The vertical axis differentiates Funders (top) from Implementers (bottom). This orthogonal arrangement creates four distinct quadrants populated by exemplars such as OCHA (Humanitarian Funder), Save the Children (Humanitarian Implementer), IADB (Development Funder) and Chemonics (Development Implementer). The analysis also highlights the strategic position of connectors like J-PAL and The Hewlett Foundation, which sit near the origin and bridge these functional zones.
@@
The separation between these clusters highlights a systemic coordination risk. The structural holes between the quadrants are bridged by relatively few actors, potentially stalling the transition from crisis response to resilience building. We identify specific connectors such as J-PAL and The Hewlett Foundation that sit near the origin of this coordinate system, suggesting they are uniquely capable of integrating these otherwise disconnected functional zones. Similar topological fissures have been observed in global health networks[17] and climate adaptation finance[18], where issue-specific clusters often fail to integrate with broader systems strengthening[19, 20].
@@
To derive this ranking, we analysed the underlying bipartite graph connecting providers to receivers. We calculated the Network Centrality (Hub Score)[21] for all organisations to measure their systemic influence. In this bipartite context, a high Hub Score identifies actors that serve as critical “hubs”, connecting to a wide and significant range of recipients across the global network.
@@
We identify a core inner ring of 25 actors who dominate the network’s structural connectivity. This visualisation reveals a critical paradox between scale and influence. While the inner ring includes the expected financial giants (US Department of State, European Commission, World Bank) it also contains a distinct class of Knowledge Brokers. These are organisations with lower financial volume but disproportionately high network centrality (Extended Data Fig. 3).
@@
Their high betweenness centrality [22] (Extended Data Fig.3) indicates both J-Pal and the Hewlett foundation sit on the shortest paths between otherwise poorly connected donor-implementer clusters. These findings overturn the assumption that influence is solely a function of budget. In the hidden geometry of global aid, smaller research-led organisations provide the essential connective tissue that holds the fragmented system together.
@@ Discussion
Structural Fragmentation. Our topological analysis reveals that the global aid system is not a cohesive or integrated network. It is rather a collection of functional and geographic islands that are sparsely connected. The clear separation observed in the UMAP embeddings between the humanitarian and development clusters points to a persistent structural divide. Organisations tend to cluster with peers that share similar operational timeframes and funding sources[23, 24]. While this specialisation allows for operational efficiency within clusters it creates systemic risk at the macro level. It enables rapid response for humanitarian actors or long-term infrastructure planning for development banks but leaves the wider system disjointed.
@@
The sparseness of connections between these clusters suggests a high potential for duplication of effort and coordination failure. The economic costs of such fragmentation are well documented, often overwhelming the bureaucratic capacity of recipient nations[25, 26] and creating game-theoretic equilibria where donors prioritise visibility over collective impact[27]. When evidence is generated in one cluster such as a randomised control trial on health interventions within the development cluster the lack of structural bridges prevents it from diffusing efficiently to the humanitarian cluster. This fragmentation explains a longstanding paradox in development economics regarding why evidence of what works often fails to scale. The network geometry confirms that the barriers to scaling are not merely financial or political[28, 24] but topological. Information simply cannot traverse the structural holes between the islands of the aid system without the intervention of specific bridging actors[29].
@@
The Strategic Role of Brokers. The identification of a solar system architecture offers a new strategic framework for donors. Traditionally aid portfolios are constructed based on volume. Funding is allocated to doers such as large NGOs and contractors capable of absorbing significant capital. However our centrality analysis demonstrates that high-volume implementers often operate within deep but narrow silos. Industry analysis of donor practices confirms that this fragmentation creates a “vicious cycle” of grant dependence where short-term reporting requirements stifle the long-term systemic growth of support organisations[30].
@@
To de-risk portfolios and accelerate systemic coordination donors must balance investments in doers with investments in connectors. Our results highlight that knowledge brokers provide disproportionate structural influence relative to their size. These are specifically academic research centers like the Abdul Latif Jameel Poverty Action Lab (J-PAL) at MIT and universities like Johns Hopkins University as well as strategic foundations like The Hewlett Foundation. These actors possess high betweenness centrality which means they sit on the shortest paths between otherwise disconnected nodes. Sociological theory posits that such weak ties are more effective for the diffusion of new information than strong ties to established partners[31]. This aligns with the rise of “knowledge actors” who exert influence through transnational standard-setting rather than financial weight[32, 33, 34]. J-PAL, for instance, leveraged this position to shift the sector toward evidence-based micro-interventions[35, 36, 37].
@@
Network Resilience. From a complex systems perspective the current architecture of global aid exhibits signs of structural fragility. The system is heavily dependent on the Donor Core. This consists of a small number of massive hubs like USAID, the World Bank and the European Commission that provide the majority of financial fuel. In network theory systems dependent on a few central hubs are resilient to random failure but highly vulnerable to targeted disruption or the abrupt removal of a hub[38, 39]. This vulnerability is not hypothetical. Recent volatility such as the recalibration of US multilateral commitments between 2017 and 2021 demonstrates how policy discontinuities in a key donor nation can rapidly propagate instability across the wider network.
Strengthening the middle layer of knowledge brokers creates a more resilient and adaptive mesh. Much like the internet’s distributed architecture allows data to reroute around damaged nodes—unlike a centralised telephone exchange which fails if the switchboard goes down, a network with a strong middle layer of brokers is less prone to fragmentation even if a major donor pulls back. These brokers act as the system’s memory and nervous system by retaining evidence and maintaining relationships that transcend political cycles. Our analysis suggests that the resilience of the global aid architecture depends on shifting from a star topology of centralised control to a mesh topology of distributed coordination[40]. The hidden geometry we identified suggests that this mesh already exists in embryonic form through the university and foundation sectors. It simply requires recognition and resource support to function as a true stabiliser. Recent systemic design scholarship reinforces this arguing that traditional management dimensions must shift toward a “complexity paradigm” to handle such emergent dynamics[41].
@@ Methods
We derived the global aid transaction network from the International Aid Transparency Initiative (IATI) registry[20], the most comprehensive open-data standard for development cooperation reporting. Transaction-level data were programmatically retrieved using the IATI API, covering the full temporal span of available records from 1967 to 2025.
The data collection pipeline iteratively queried the IATI transactions endpoint with pagination to ensure complete coverage across all reporting organisations. Each API response was parsed and stored in a structured format, retaining core fields including transaction identifiers, reporting organisation, provider organisation, receiver organisation, transaction date, transaction value, currency, and transaction type. To ensure consistency across reporting practices, all transaction values were filtered to exclude null or zero-value entries, and only valid financial flows were retained for analysis. Organisation names in the raw IATI records exhibit substantial heterogeneity due to inconsistent reporting conventions, abbreviations, and spelling variants. To address this, we applied a systematic normalisation procedure combining exact matching and fuzzy string matching to resolve common aliases (e.g., mapping “USAID” to “United States Agency for International Development”). This process reduced duplication and ensured that each organisation was represented as a unique node in the network.
Transactions were then aggregated at the organisation–organisation level, linking provider and receiver entities through observed financial flows. Records with missing or unresolvable organisation identifiers were excluded. The final cleaned dataset comprised over 10 million transactions involving 2,456 distinct publishing organisations operating across 230 countries and territories. All preprocessing steps were implemented in Python to ensure reproducibility. The resulting cleaned and normalised transaction table served as the basis for subsequent network construction, projection, and analysis.
@@
We modelled the global aid ecosystem as a bipartite (two-mode) network derived from the cleaned IATI transaction dataset. In this representation, nodes correspond to organisations, and edges represent observed financial transactions between a provider organisation and a receiver organisation[42, 43, 44]. Each transaction induces a directed edge from provider to receiver, reflecting the flow of resources within the aid system.
To construct the network, we first grouped transactions by unique provider–receiver organisation pairs. Multiple transactions between the same pair were aggregated into a single weighted edge, with the edge weight corresponding to transaction frequency rather than monetary value. This choice emphasises sustained relational activity and repeated collaboration over raw financial volume, allowing the network topology to capture structural connectivity rather than scale alone.
The resulting bipartite graph includes all organisations that appear at least once as either a provider or a receiver in the transaction corpus. Self-loops and records with unresolved organisation identifiers were excluded. The graph was implemented using the NetworkX library, enabling efficient handling of large-scale relational data and the computation of downstream structural metrics.
@@
To reveal latent structural patterns in the global aid network, we applied a two-stage representation-learning and dimensionality-reduction pipeline based on node embeddings, followed by a nonlinear projection. This approach enables the comparison of organisations based on their topological roles rather than their geographic location or financial volume.
We first generated high-dimensional vector representations of organisations using the node2vec algorithm[45] applied to the bipartite graph. Node2vec performs biased random walks over the network to learn embeddings that capture both local neighbourhood structure and broader structural equivalence. Walks were generated from each node and optimised using a skip-gram objective, producing fixed-length embeddings for all organisations in the network. These embeddings encode similarities in funding and implementation patterns, allowing organisations with comparable network positions to be mapped close together in embedding space, even if they are not directly connected. The resulting high-dimensional embeddings were then projected into two dimensions using Uniform Manifold Approximation and Projection (UMAP)[46]. UMAP is a nonlinear dimensionality reduction technique that preserves local neighbourhood relationships while maintaining meaningful global structure. We used a cosine distance metric on the embedding vectors, with standard parameter settings (including a neighbourhood size of 15 and a minimum distance of 0.1), to obtain a stable two-dimensional manifold representation of the network.
The 2D projection reveals a clear geometric organisation of the aid ecosystem, in which organisations cluster by functional role and patterns of interaction. Rather than imposing predefined categories, clusters emerge organically from similarities in network connectivity. These clusters correspond to systematic differences between humanitarian and development actors, as well as between funding-oriented and implementation-oriented organisations. This embedding-based approach enables the identification of structural proximity and separation within the aid system without reliance on exogenous labels. The resulting low-dimensional representation forms the basis for the visual analyses presented in the Results section. It supports the identification of structural bridges and coordination gaps across the global aid network.
@@
We calculated network centrality metrics to identify the core architecture of the system. Degree Centrality was used to measure the total connectivity of actors. Betweenness Centrality was calculated to quantify brokerage[47]. This identifies actors that lie on the shortest paths between other pairs of organisations[48]. We validated these offline network metrics by correlating them with online visibility metrics using network centrality measures of linkage patterns from the organisations’ websites. We found a strong positive Pearson correlation of 0.48 between the structural position of an organisation in the IATI network and the centrality of its website in the World Wide Web (Global PageRank) (Extended Data Fig. 4). The solar system visualisation ranks the top 100 organisations by a composite centrality score with the inner ring representing the top 25 actors. To handle the scale of the graph we employed advanced approximation algorithms for neighbourhood functions[49]. This methodological approach aligns with previous efforts to map aid flows[50, 51, 52] but extends the analysis to a global scale, allowing for the detection of critical transitions[53] and systemic risks[54, 55].
@@
The dataset analysed during the current study is publicly available in the International Aid Transparency Initiative (IATI) Registry. The raw XML data can be accessed at IATI API https://api.iatistandard.org/datastore. The processed network files and edge lists used for the topological analysis are available in the IATI_Global_Aid repository at https://github.com/leagueofscholars/IATI_Global_Aid.git.
@@ Acknowledgements
We thank the International Aid Transparency Initiative (IATI) for maintaining and providing open access to the global aid data that made this research possible. We also acknowledge Snowmelt for its support and for valuable discussions that helped inform the broader context of this work.
@@
P.X.M. conceived the study, developed the core ideas, and led the writing of the manuscript. X.G. collected and curated the data and contributed to the empirical analysis. M.-A.R. and P.B. provided conceptual guidance, methodological insights, and critical feedback that shaped the analysis and interpretation. All authors reviewed and approved the final manuscript.
Additional Information
The author(s) declare no competing interests.